# Inverse problem for semidiscrete stochastic parabolic operators in arbitrary dimensions

### v1

@@ 1. Introduction
The uniqueness follows directly from Theorem 1.1: if $\Lambda_{1}(g_{1})=\Lambda_{1}(g_{2})$ $\mathbb{P}$-a.s., then $g_{1}=g_{2}$ in $Q$, $\mathbb{P}$-a.s.
Semidiscrete Cauchy Problem. We now consider the same semi-discretization of (1.1), but with $g=0$ and a non-homogeneous Dirichlet boundary condition:
[…]
Although the discretization is structurally similar to the previous problem, the key distinction lies in the type of data: Here, we observe the boundary values and aim to recover the interior solution.
We define the observation operator $\Lambda_{2}:L^{2}_{\mathcal{F}}(0,T;H_{h}^{2}(\mathcal{M}))\to X_{2}$ by
$\left.\Lambda_{2}(w)\triangleq\left(w,\partial_{\nu}w\right)\right|_{\Gamma^{+}\times(0,T)},$
where $w$ solves (1.8). In this case, for the space $X_{2}\triangleq H^{1}_{\mathcal{F}}(0,T;H_{h}^{1}(\partial\mathcal{M})\cap H_{h}^{1/2}(\partial\mathcal{M}))\times\left(L^{2}_{\mathcal{F}}(0,T;L_{h}^{2}(\Gamma^{+}))\right)^{n}$ adopted the norm defined for
$\left\|\left(f_{1},f_{2},...,f_{n},f_{n+1}\right)\right\|_{X_{2}}\triangleq\|f_{1}\|_{H^{1}_{\mathcal{F}}(0,T;H_{h}^{1}(\partial\mathcal{M})\cap H_{h}^{1/2}(\partial\mathcal{M}))}+\sum_{i=2}^{n+1}\|f_{i}\|_{L^{2}_{\mathcal{F}}(0,T;L_{h}^{2}(\Gamma^{+}))},$
for $f_{1}\in H^{1}_{\mathcal{F}}(0,T;H_{h}^{1}(\partial\mathcal{M})\cap H_{h}^{1/2}(\partial\mathcal{M}))$ and $f_{i}\in L^{2}_{\mathcal{F}}(0,T;L_{h}^{2}(\Gamma^{+}))$ with $i=2\cdots n+1$.
The inverse problem consists of determining the solution $w$ in a subdomain $\mathcal{M}_{0}\subset\mathcal{M}$ over the time interval $(\varepsilon,T-\varepsilon)$, based on the boundary measurement given by $\Lambda_{2}(w)$. We establish the following Hölder-type stability estimate.
@@
Let $a_{1i},a_{2}\in L^{\infty}_{\mathcal{F}}(0,T;L_{h}^{\infty}(\mathcal{M}))$, $a_{3}\in L^{\infty}_{\mathcal{F}}(0,T;W_{h}^{1,\infty}(\mathcal{M}))$, $\xi\in H^{1}_{\mathcal{F}}(0,T;H_{h}^{1}(\partial\mathcal{M})\cap H_{h}^{1/2}(\partial\mathcal{M)})$, and let $\varepsilon>0$. Then there exist constants $C>0$, $h^{*}>0$, and $\kappa\in(0,1)$ such that for all $h\in(0,h^{*})$ and any solution $w\in L^{2}_{\mathcal{F}}(0,T;H_{h}^{2}(\mathcal{M}))$ of (1.8) satisfying
@@
(1.9) $\|w\|_{L^{2}_{\mathcal{F}}(\varepsilon,T-\varepsilon;H_{h}^{2}(\mathcal{M}_{0}))}\leq C\max{\left\{\|\Lambda_{2}(w)\|_{X_{2}},\mathbf{M}e^{-Ch^{-1}},\mathbf{M}^{\kappa}\|\Lambda_{2}(w)\|_{X_{2}}^{1-\kappa}\right\}}.$
@@
The size of the observation set $\Gamma^{+}$ depends on the choice of the reference point $x^{*}$. If $x^{*}$ is chosen so that the vector $x-x^{*}$ is perpendicular to the face of $\partial\mathcal{M}$, then $\Gamma^{+}$ is smaller. In contrast, if $x^{*}$ is chosen in a more oblique direction, $\Gamma^{+}$ increases in size. Thus, an appropriate choice of $x^{*}$ could minimize the required observation region.
Remark 1.5.
We observe that if we consider $h$ to be small enough and the norm $|\Lambda_{2}(w)|{X_{2}}$ to be small compared to the constant $\mathbf{M}$ (i.e., $|\Lambda_{2}(w)|{X_{2}}<\mathbf{M}$), we can recover the estimate presented in WWW:2024 . In that reference, the first two terms in (1.9) were neglected.
@@
The two inverse problems under consideration are associated with stochastic parabolic equations. Stochastic partial differential equations (SPDEs) have broad applications across multiple disciplines. They model phenomena such as particle diffusion and heat propagation in physics, option pricing in finance, epidemic spread and population dynamics in biology, systems with uncertain parameters in engineering, and pollutant dispersion in environmental sciences. These equations incorporate randomness and uncertainty, making them particularly suitable for modeling inherently unpredictable phenomena. Consequently, stochastic PDEs often provide more realistic representations than their deterministic counterparts. Nevertheless, research on inverse problems for stochastic PDEs remains limited. For further details, see (lu2021mathematical, , Chapter 5) and the references therein.
@@
For the inverse source problem related to stochastic parabolic equations, ref27 establishes uniqueness in determining the source function $f$ in the drift term. The inverse source problem of simultaneously determining two types of sources, $f$ (drift term) and $g$ (diffusion term), in a stochastic parabolic equation was studied in ref37 using observations at the final time and on the lateral boundary. A unique result for this inverse problem was obtained using a global Carleman estimate for the continuous stochastic parabolic equation. In addition, ref1 and ref2 discuss the application of regularization techniques within numerical methods to inverse random source problems.
@@
The Cauchy problem aims to recover the solution using the observed data on the lateral boundary. In Yamamoto_2009 , conditional stability was established for the Cauchy problem of deterministic parabolic equations. The conditional stability and the convergence rate of the Tikhonov regularization method for the Cauchy problem of stochastic parabolic equations were obtained in ref13 . It is worth mentioning that ref28 provides a detailed review of inverse problems for stochastic PDEs. However, all of these results were obtained within a continuous framework. For semidiscrete inverse problems related to stochastic differential equations, see WWW:2024 , which studies the one-dimensional case of both issues, including a semidiscrete inverse random source problem and a semidiscrete Cauchy problem.
@@
This article builds upon the results of WWW:2024 , extending them to arbitrary spatial dimensions and providing several refinements. In particular, Theorem 1.1 provides integrability conditions for the coefficient $a_{2}$ that depend on the spatial dimension of the problem. Furthermore, Theorem 1.3 establishes a stability estimate in which the right-hand side of inequality (1.9) explicitly reveals the dependence on the mesh size $h$ and the size of the observation operator. As a consequence, the second result recovers the same estimate as in WWW:2024 when $h$ and the observation region are chosen sufficiently small.
The main tools to address the previous inverse problems are two Carleman estimates. The first is for a homogeneous Dirichlet condition.
@@
Let $\mbox{reg}^{0}>0$, $\psi$ satisfy assumptions (2.1) and $\varphi$ according to (2.2). For $\lambda_{1}\geq 1$ sufficiently large, there exist $C$, $\tau_{0}\geq 1$, $h_{0}>0$, $\varepsilon_{0}>0$, depending on $G$, $\beta$, $\mbox{reg}^{0}$, $T$, and $\lambda$, such that for any $\gamma$, with $\mbox{reg}(\gamma)\leq\mbox{reg}^{0}$ we have
(1.10) $\begin{split}&J(w)+\mathbb{E}\int_{Q}s^{3}\lambda^{4}_{1}\varphi^{3}\,e^{2s\varphi}|w|^{2}\,dt+\sum_{i,j=1}^{n}\mathbb{E}\int_{Q_{ij}^{\ast}}s^{-1}e^{2s\varphi}|D_{ij}^{2}w|^{2}\,dt+\mathbb{E}\int_{Q}s\lambda_{1}^{2}\varphi\,e^{2s\varphi}|g|^{2}dt\\ &\leq\,C_{s_{0},\varepsilon}\left(\mathbb{E}\int_{Q}\,e^{2s\varphi}|f|^{2}\,dt+\sum_{i=1}^{n}\mathbb{E}\int_{Q_{i}^{\ast}}\,e^{2s\varphi}|D_{i}g|^{2}dt+\sum_{i=1}^{n}\mathbb{E}\int_{0}^{T}\int_{\Gamma^{+}}st_{r}^{i}(e^{2s\varphi}|D_{i}w|^{2})\,dt\right.\\ &\left.+\left.\mathbb{E}\int_{\mathcal{M}}s^{2}\,e^{2s\varphi}|w|^{2}\right|_{t=T}\right).\end{split}$
for all $\tau\geq\tau_{0}$, $0<h\leq h_{0}$, $s(t)h\leq\epsilon_{0}$, where $J(w)\triangleq\sum_{i=1}^{n}\mathbb{E}\int_{Q_{i}^{\ast}}s\lambda_{1}^{2}\varphi e^{2s\theta\varphi}\,|D_{i}w|^{2}\,dt+\mathbb{E}\int_{Q}s\lambda_{1}^{2}\varphi e^{2s\theta\varphi}|A_{i}D_{i}w|^{2}\,dt$, $f\in L^{2}_{\mathcal{F}}(0,T;L^{2}(\mathcal{M}))$, $g\in L^{2}_{\mathcal{F}}(0,T;H^{1}(\mathcal{M}))$, and $w$ satisfies $dw-\sum_{i=1}^{n}D_{i}(\gamma_{i}D_{i}w)\,dt=fdt+gdB(t)$ with $w=0$ on $\partial\mathcal{M}$ and $w(0)=0$ in $\mathcal{M}$.
The second is for the nonhomogeneous Dirichlet condition given in Theorem 3.3. Theorem 1.6 and Theorem 3.3 extend the one-dimensional Carleman estimates presented in WWW:2024 . We adopt the weight function introduced in that work. However, in the multidimensional setting, additional difficulties arise due to the need to incorporate second-order terms $D^{2}_{ij}$ and the fact that the boundary is no longer a simple pair of points. To establish these estimates, we rely on fundamental discrete tools, such as integration by parts and the product rule, which we state below.
Proposition 1.7 ((LOPD, , Lemma 2.2)).
@@
Proposition 1.8 ((boyer-2010-elliptic, , Lemma 2.1 and 2.2)).
@@
The remainder of this paper is organized as follows. In Section 2, we prove a Lipschitz stability result of the space-semidiscrete inverse source problem as a direct result of the Carleman estimate presented in the Appendix B. Finally, in Section 3, we study the space-semidiscrete inverse Cauchy problem and establish a Hölder stability result based on a second Carleman estimate and a geometric argument.
2. Semidiscrete inverse random source problem
In this section, we establish the stability of the discrete inverse random source problem (Theorem 1.1). The key tool for proving this result is a semidiscrete Carleman estimate, whose proof is deferred to B. To formulate this estimate, we first introduce the necessary assumptions on $\gamma\triangleq(\gamma_{1},...,\gamma_{n})$ and the associated weight functions. Assume $\gamma_{i}>0$ for all i=1,,ni=1,...,n with
$\mbox{reg}(\gamma)\triangleq\text{ess}\sup_{\begin{subarray}{c}x\in G\\ i=1,\cdots,n\end{subarray}}\left(\gamma_{i}+\frac{1}{\gamma_{i}}+\sum_{i=1}^{n}|\partial_{i}\gamma_{i}|^{2}\right)<\infty.$
Now, let $x^{\ast}\notin\overline{G}$ and, for $\lambda>1$, define the weight functions
(2.1) $\psi(x)$ $=|x-x^{*}|^{2},$
(2.2) $\varphi(t,x)$ $=e^{\lambda\psi(x)},$
and for $t_{0}\in(0,T)$ and $\beta>0$,

### v2

@@ 1. Introduction
Here we provide an example that satisfies (1.6). Indeed,
$g(x)=\prod_{k=1}^{n}(x_{k}+C^{\prime})\mathcal{S}(t,\omega),$
for any constant $C^{\prime}>0$ and $\mathcal{S}\in L^{2}_{\mathcal{F}}(0,T);\mathbb{R})$ is a real-valued adapted process satisfying $\mathcal{S}(t,\omega)>0$ -$\mathbb{P}\text{-a.s.}$. To verify condition (1.6), fix a direction $i\in\{1,\ldots,n\}$ and write $g(x)=(x_{i}+C^{\prime})\cdot R_{i}(x)$, where
$R_{i}(x):=\prod_{k\neq i}(x_{k}+C^{\prime})\mathcal{S}(t,\omega).$
Since the discrete operators $D_{i}$ and $A_{i}$ act only on the $i$-th variable, a direct computation on $\mathcal{M}_{i}^{\ast}$ gives
$D_{i}g(x)=\frac{g\!\left(x+\tfrac{h}{2}e_{i}\right)-g\!\left(x-\tfrac{h}{2}e_{i}\right)}{h}=\frac{\bigl(x_{i}+\tfrac{h}{2}+C^{\prime}\bigr)-\bigl(x_{i}-\tfrac{h}{2}+C^{\prime}\bigr)}{h}\cdot R_{i}(x)=R_{i}(x),$
$A_{i}g(x)=\frac{g\!\left(x+\tfrac{h}{2}e_{i}\right)+g\!\left(x-\tfrac{h}{2}e_{i}\right)}{2}=(x_{i}+C^{\prime})\cdot R_{i}(x).$
Therefore,
$\frac{|D_{i}g(x)|}{|A_{i}g(x)|}=\frac{R_{i}(x)}{(x_{i}+C^{\prime})\,R_{i}(x)}=\frac{1}{x_{i}+C^{\prime}}\leq\frac{1}{C^{\prime}},\qquad\forall\,x\in\mathcal{M}_{i}^{\ast},$
where we used $x_{i}\geq 0$ and $\mathcal{S}(t,\omega)>0$ -$\mathbb{P}\text{-a.s.}$. Hence condition (1.6) holds with constant $C=1/C^{\prime}$, uniformly in $h$, for every $i=1,\ldots,n$. More generally, it is possible to build a family of functions that verify that condition. For any adapted process $g\in L^{2}_{\mathcal{F}}(0,T;H_{h}^{1}(\mathcal{M}))$ satisfying $\|D_{i}g\|_{L^{\infty}_{\mathcal{F}}(0,T;L^{\infty}_{h}(\mathcal{M}_{i}))}<C^{\prime}$ and $|A_{i}g(x,t,\omega)|\geq\alpha>0$ for all $x\in\mathcal{M}_{i}$, $t\in(0,T)$, -$\mathbb{P}\text{-a.s.}$; it follows that $|D_{i}g(x,t,\omega)|\leq C^{\prime}=\frac{C^{\prime}}{\alpha}\cdot\alpha\leq C|A_{i}g(x,t,\omega)|$ with $C=\tfrac{C^{\prime}}{\alpha}$.
Semi-discrete Cauchy Problem. We consider the same spatial semi-discretization of (1.1), with $g=0$:
(1.8) $\mathcal{A}_{h}w=\left(\sum_{i=1}^{n}a_{1i}A_{i}D_{i}(w)+a_{2}w\right)dt+a_{3}w\,dB(t),\quad\text{in }Q,$
While the discretization shares structural similarities with the previous inverse problem, the nature of the available data differs. In this setting, our objective is to recover $w$ within a subdomain $D\subset Q$. This is achieved through the observation operator $\Lambda_{2}:L^{2}_{\mathcal{F}}(0,T;H_{h}^{2}(\mathcal{M}))\to X_{2}$, which is defined as follows:
$\Lambda_{2}(w):=\left(\left.w\right|_{\Gamma\times(0,T)},\left.\partial_{\nu}w\right|_{\Gamma\times(0,T)}\right),$
where $w$ solves (1.8). The normal derivative is given by
$\partial_{\nu}w:=\sum_{i=1}^{n}t_{r}^{i}(D_{i}w)e_{i}\nu_{i},$
and the set $\Gamma=\Gamma_{0}\cap\partial M$, where $\Gamma_{0}$ is an arbitrary non-empty open subset of $\partial G$. In this case, we set $X_{2}:=H^{1}_{\mathcal{F}}(0,T;H_{h}^{1}(\Gamma)\cap H_{h}^{1/2}(\Gamma))\times\left(L^{2}_{\mathcal{F}}(0,T;L_{h}^{2}(\Gamma))\right)^{n}$ endorsed with norm
$\left\|\left(f_{1},f_{2},...,f_{n},f_{n+1}\right)\right\|_{X_{2}}:=\|f_{1}\|_{H^{1}_{\mathcal{F}}(0,T;H_{h}^{1}(\Gamma)\cap H_{h}^{1/2}(\Gamma))}+\sum_{i=2}^{n+1}\|f_{i}\|_{L^{2}_{\mathcal{F}}(0,T;L_{h}^{2}(\Gamma))},$
for $f_{1}\in H^{1}_{\mathcal{F}}(0,T;H_{h}^{1}(\Gamma)\cap H_{h}^{1/2}(\Gamma))$ and $f_{i}\in L^{2}_{\mathcal{F}}(0,T;L_{h}^{2}(\Gamma))$ with $i=2,\ldots,n+1$.
The inverse problem consists of determining the solution $w$ in an open subdomain $\mathcal{M}_{0}\subset\mathcal{M}$ such that $\overline{\mathcal{M}}_{0}\subset\mathcal{M}\cup\Gamma$ and $\partial\mathcal{M}_{0}\cap\partial\mathcal{M}\subsetneq\Gamma$ over the time interval $(\varepsilon,T-\varepsilon)$, based on the boundary measurement given by $\Lambda_{2}(w)$. We establish the following Hölder-type stability estimate:
@@
Let $a_{1i},a_{2}\in L^{\infty}_{\mathcal{F}}(0,T;L_{h}^{\infty}(\mathcal{M}))$, $a_{3}\in L^{\infty}_{\mathcal{F}}(0,T;W_{h}^{1,\infty}(\mathcal{M}))$, $\xi\in H^{1}_{\mathcal{F}}(0,T;H_{h}^{1}(\Gamma)\cap H_{h}^{1/2}(\Gamma))$, and let $\varepsilon>0$. Then there exist constants $\mathcal{C}>0$, $h^{*}>0$, and $\kappa\in(0,1)$ such that for all $h\in(0,h^{*})$ and any solution $w\in L^{2}_{\mathcal{F}}(0,T;H_{h}^{2}(\mathcal{M}))$ of (1.8) with $w=\xi\;\text{on}\;\,\Gamma\times(0,T)$ satisfying
@@
(1.9) $\|w\|_{L^{2}_{\mathcal{F}}(\varepsilon,T-\varepsilon;H_{h}^{2}(\mathcal{M}_{0}))}\leq\mathcal{C}\max{\left\{\|\Lambda_{2}(w)\|_{X_{2}},\mathbf{M}e^{-\mathcal{C}h^{-1}},\mathbf{M}^{\kappa}\|\Lambda_{2}(w)\|_{X_{2}}^{1-\kappa}\right\}}.$
@@
We observe that by refining the mesh size $h$ and the observation region such that $\|\Lambda_{2}(w)\|_{X_{2}}<\mathbf{M}$, inequality (1.9) can be rewritten as
(1.10) $\|w\|_{L^{2}_{\mathcal{F}}(\varepsilon,T-\varepsilon;H_{h}^{2}(\mathcal{M}_{0}))}\leq\mathcal{C}\mathbf{M}^{\kappa}\|\Lambda_{2}(w)\|_{X_{2}}^{1-\kappa}.$
The above inequality stands for a generalization of (Wu et al., 2024, Theorem 2.6). Moreover, in (Wu et al., 2024, Remark 2.11) the authors claim that (Wu et al., 2024, Theorem 2.6) implies the uniqueness of this inverse problem. This latter conclusion is incorrect. Indeed, by examining the proof of that Theorem and assuming that the measurements are zero, we see that in the unlabeled inequality preceding (Wu et al., 2024, Inequality 5.16), an exponential error term remains, which prevents one to concluding a uniqueness result. For this reason, in inequality (1.9) appears the term $\mathbf{M}e^{-\mathcal{C}h^{-1}}$. There is another way to see that a uniqueness result is not possible. From (1.10), we obtain, up to a constant, $\mathcal{C}e^{-h^{-1}}\leq\|\Lambda_{2}(w)\|^{1-\kappa}_{X_{2}}$. Then, assuming zero measurement does not imply uniqueness in this inverse problem. This phenomenon is not new in discrete inverse problem: in the continuous case, that term vanishes when the Carleman parameter is taken sufficiently large, while in the discrete setting this is not allowed because the discrete mesh size and the Carleman parameter are connected, see Theorems 1.5 and 3.3. Moreover, similar results are obtained in several estimates for discrete systems; see for instance Ervedoza and de Gournay (2011); Lecaros et al. (2023).
@@
The two inverse problems under consideration are associated with stochastic parabolic equations. Stochastic partial differential equations (SPDEs) have broad applications across multiple disciplines. They model phenomena such as particle diffusion and heat propagation in physics, option pricing in finance, epidemic spread and population dynamics in biology, systems with uncertain parameters in engineering, and pollutant dispersion in environmental sciences. These equations incorporate randomness and uncertainty, making them particularly suitable for modeling inherently unpredictable phenomena. Consequently, stochastic PDEs often provide more realistic representations than their deterministic counterparts. Nevertheless, research on inverse problems for stochastic PDEs remains limited. For further details, see (Lü and Zhang, 2021, Chapter 5) and the references therein.
@@
For the inverse source problem related to stochastic parabolic equations, Lü (2012) establishes uniqueness in determining the source function $f$ in the drift term. The inverse source problem of simultaneously determining two types of sources, $f$ (drift term) and $g$ (diffusion term), in a stochastic parabolic equation was studied in Yuan (2021) using observations at the final time and on the lateral boundary. A unique result for this inverse problem was obtained using a global Carleman estimate for the continuous stochastic parabolic equation. In addition, Bao et al. (2010) and Bao and Xu (2012) discuss the application of regularization techniques within numerical methods to inverse random source problems.
@@
The Cauchy problem aims to recover the solution using the observed data on the lateral boundary. In Yamamoto (2009), conditional stability was established for the Cauchy problem of deterministic parabolic equations. The conditional stability and the convergence rate of the Tikhonov regularization method for the Cauchy problem of stochastic parabolic equations were obtained in Dou et al. (2024). However, all of the above cited results were obtained within a continuous framework, as is the survey Lü and Zhang (2024). For semi-discrete inverse problems related to stochastic differential equations, see Wu et al. (2024), which studies the one-dimensional case of both issues, including a semi-discrete inverse random source problem and a semi-discrete Cauchy problem.
@@
This article builds upon the results of Wu et al. (2024), extending them to arbitrary spatial dimensions and providing several refinements. In particular, Theorem 1.1 provides integrability conditions for the coefficient $a_{2}$ that depend on the spatial dimension of the problem. Furthermore, Theorem 1.3 establishes a stability estimate in which the right-hand side of inequality (1.9) explicitly reveals the dependence on the mesh size $h$ and the size of the observation operator. As a consequence, the second result recovers the same estimate as in Wu et al. (2024) when $h$ and the observation region are chosen sufficiently small. The main tools to address the previous inverse problems are two Carleman estimates. The first is for a homogeneous Dirichlet condition.
@@
Let $\psi$ satisfying assumptions 2.1 and $\varphi=e^{\lambda\psi}$. For $\lambda_{1}\geq 1$ sufficiently large, there exist $\mathcal{C}$, $\tau_{0}\geq 1$, $h_{0}>0$, $\varepsilon_{0}>0$, depending on $G,\,G_{0}$, $\beta$, $c_{0}$, $T$, and $\lambda$, such that
(1.11) $\begin{split}&J(w)+\mathbb{E}\int_{Q}s^{3}\lambda^{4}_{1}\varphi^{3}\,e^{2s\varphi}|w|^{2}\,dt+\sum_{i,j=1}^{n}\mathbb{E}\int_{Q_{ij}^{\ast}}s^{-1}e^{2s\varphi}|D_{ij}^{2}w|^{2}\,dt+\mathbb{E}\int_{Q}s\lambda_{1}^{2}\varphi\,e^{2s\varphi}|g|^{2}dt\\ &\leq\,\mathcal{C}_{\tau_{0},\varepsilon_{0}}\left(\mathbb{E}\int_{Q}\,e^{2s\varphi}|f|^{2}\,dt+\sum_{i=1}^{n}\mathbb{E}\int_{Q_{i}^{\ast}}\,e^{2s\varphi}|D_{i}g|^{2}dt+\int_{0}^{T}\int_{G_{0}\cap\mathcal{M}}s^{3}\varphi^{3}e^{2s\varphi}|w|^{2}\,dt\right.\\ &\left.+\left.\mathbb{E}\int_{\mathcal{M}}s^{2}\,e^{2s\varphi}|w|^{2}\right|_{t=T}\right),\end{split}$
for all $\tau\geq\tau_{0}$, $0<h\leq h_{0}$, $s(t)h\leq\epsilon_{0}$, where $J(w):=\sum_{i=1}^{n}\mathbb{E}\int_{Q_{i}^{\ast}}s\lambda_{1}^{2}\varphi e^{2s\varphi}\,|D_{i}w|^{2}\,dt+\mathbb{E}\int_{Q}s\lambda_{1}^{2}\varphi e^{2s\varphi}|A_{i}D_{i}w|^{2}\,dt$, $f\in L^{2}_{\mathcal{F}}(0,T;L^{2}_{h}(\mathcal{M}))$, $g\in L^{2}_{\mathcal{F}}(0,T;H^{1}_{h}(\mathcal{M}))$, and $w$ satisfies $dw-\sum_{i=1}^{n}D_{i}(\gamma_{i}D_{i}w)\,dt=fdt+gdB(t)$ with $w=0$ on $\partial\mathcal{M}$ and $w|_{t=0}=0$ in $\mathcal{M}$.
The second is for the nonhomogeneous Dirichlet condition given in Theorem 3.7. Theorem 3.3 provides, to our knowledge, the first Carleman estimate for a stochastic forward semi-discrete parabolic operator with interior observation. Moreover, Theorems 1.53.7 extend the one-dimensional Carleman estimates established in Wu et al. (2024). We adopt the general structure of the weight function introduced in that work, but modify its spatial component in order to treat more general configurations. This modification allows us to derive Carleman estimates corresponding to interior or boundary observations on arbitrary subsets of the corresponding domains. However, in the multidimensional setting, additional difficulties arise due to the needed to incorporate second-order terms $D^{2}_{ij}$ and the fact that the boundary is no longer a simple pair of points. To establish these estimates, we rely on fundamental discrete tools, such as integration by parts and the product rule, which we state below.
Proposition 1.6 ((Lecaros et al., 2023, Lemma 2.2)).
@@
Proposition 1.7 ((Boyer et al., 2010b, Lemma 2.1 and 2.2)).
@@
The remainder of this paper is organized as follows. In Section 2, we prove a Lipschitz stability result of the spatial semi-discrete inverse random source problem applying the result of the Carleman estimate (1.11) where its proof is presented in the Appendix A. Finally, in Section 3, we study the spatial semi-discrete inverse Cauchy problem and establish a Hölder stability result based on a second Carleman estimate and a geometric argument.
2. Semi-discrete inverse random source problem
In this section, we establish the stability of the semi-discrete inverse random source problem (Theorem 1.1). The key tool for proving this result is a semi-discrete Carleman estimate, whose proof is deferred to Appendix A. To formulate this estimate, the weight function is defined as $r=e^{s\varphi}$ with $\varphi=e^{\lambda\psi(x)}$, for $\lambda>1$ and with $\psi$ satisfies the following assumption. The construction of this weight function is classic (see e.g Èmanuilov (1995)) and the existence of this function in the semi-discrete setting is given in (Boyer et al., 2010b, Appendix A).
Assumption 2.1.
[…]
$\begin{array}[]{ccc}\psi>0\;\text{in}\;\tilde{G},\;\psi=0\;\text{on}\;\partial\tilde{G},&|\nabla\psi|\geq c\;\;\text{in}\;\tilde{G}\setminus\tilde{G}_{0},&\text{and}\;\partial_{\nu_{i}}\psi(x)\leq-c<0,\;\text{for}\,\;x\in V_{\partial_{i}G},\end{array}$
where $V_{\partial_{i}G}$ is a suffciently small neighborhood of $\partial_{i}G$ in $\tilde{G}$, in which the outward unit normal $\nu_{i}$ to $G$ is extended from $\partial_{i}G$.
On the other hand, for the temporal weight function, we define it as: for $t_{0}\in(0,T)$ and $\beta>0$,